# Mapping surface height dynamics to subsurface flow physics in free-surface turbulent flow using a shallow recurrent decoder

### v1

@@ Appendix A Parametric study of optimal rank value
To quantify error as a function of rank number, the selected error metrics are averaged depthwise, and calculated for a range of rank truncation numbers. Fig. 13 shows the rank-dependency of the SHRED error metrics for all four cases. The metrics are normalized by their largest value for scaling and comparison purposes. The chosen rank truncation number for all of the SHRED analysis in Section IV, are marked by the dashed black lines. The range of the rank analysis is different between simulation and experimental cases, due to the difference in total number of modes (equal to number of time steps). For example, the full rank of the experiments is . while for the DNS cases it is for S1 and for S2.

### v2

@@ Appendix A Parametric study of optimal rank value
To quantify error as a function of rank number, the selected error metrics are averaged depthwise, and calculated for a range of rank truncation numbers. Fig. 13 shows the rank-dependency of the SHRED error metrics for all four flow cases. The metrics are normalized by their largest value for scaling and comparison purposes. The chosen rank truncation number for all of the SHRED analysis in Section IV, are marked by the dashed black lines. The range of the rank analysis is different between simulation and experimental datasets, due to the difference in total number of modes (equal to number of time steps). For example, the full rank of the experiments is . while for the DNS cases it is for S1 and for S2.